OpenStudy (anonymous):

-5/6x =-7/8 solve using the multiplacation principle

OpenStudy (anonymous):

flip and time

OpenStudy (anonymous):

Divide -&/8 by 5/6 and you will get your answer

OpenStudy (anonymous):

Your answer is -1.05 for x

OpenStudy (anonymous):

divide both sides by -5/6 is the way i'd do -7/8 / -5/6 = 42/40 = 21/10 = 2 .1

OpenStudy (anonymous):

sorry 21/20 = 1.05

OpenStudy (anonymous):

(-5*8)=(6x*-7) -40 = -42x -40\-42 =x -0.952 = x

OpenStudy (anonymous):

oh - we all get different answers

OpenStudy (anonymous):

panther u have a key for this question?

OpenStudy (anonymous):

x = 7/8 * 6/5 = 42/40 = 21/20

OpenStudy (anonymous):

no i dont

OpenStudy (anonymous):

x must be 21/20, you can check by plugging it back in to (-5/6) * (21/20) whhich should give u -7/8.

OpenStudy (anonymous):

mine is correct

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

all it tells me is it has to be in integer or fraction form

OpenStudy (anonymous):

to divide -7/8 by -5/6 you turn -5/6 over to give -6/5 then you multiply -7/8 X -6/5 = 42/40

OpenStudy (anonymous):

fraction than x = 40\41 x = 20\21

OpenStudy (anonymous):

we have different answer, u may repost tomorrow to see what answer from some one

OpenStudy (anonymous):

in fraction form the answer is 21/20 or 1 1/20

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Thank for medal

OpenStudy (anonymous):

yw